A standard drug used to relieve pain after a certain type of operation has been successful with 80% of patients. A new drug is tried, in the belief that it will be more successful, on 200 patients, and 170 of these gain relief. Is the belief in the new drug justified? What would be said if (i) 152 or (ii) 140 of the patients had gained relief?
1
Expert's answer
2020-04-20T15:27:03-0400

200 patients-100%

170 patients-x%

"x=\frac{170\cdot100}{200}=85 \%"

85%>80%

The belief in the new drug is justified

i)

200 patients-100%

152 patients-x%

"x=\frac{152\cdot100}{200}=76 \%"

76%<80%

The belief in the new drug is not justified

ii)

200 patients-100%

140 patients-x%

"x=\frac{140\cdot100}{200}=70 \%"

70%<80%

The belief in the new drug is not justified